# Analysis of Internet of Things implementation barriers in the cold supply chain: an integrated ISM-MICMAC and DEMATEL approach

### v2

@@ 3.2.2 Step 2: Developing reachability matrix
The SSIM generates a reachability matrix; four symbols are converted to 0s and 1s to implement SSIM in the initial reachability matrix. matrix.
@@ 3.2.4 Step 4: Developing level partition
Level partition refers to categorizing or classifying distinct levels inside the information system. This can involve dividing the system into levels or hierarchies to understand its structure at various levels better. The resultant final reachability matrix is segmented into several hierarchical levels.
@@ 3.2.5 Step 5: Developing conical form
The conical form describes structuring or organizing the information system in an arrangement that resembles a conical shape, indicating a hierarchical structure. The conical form of the reachability matrix is subsequently created.
@@ 3.2.6 Step 6: Developing ISM model
A directed graph is constructed using a reachability matrix, eliminating transitive linkages. By exchanging the nodes of variables for statements, it is then converted into an ISM from the resultant digraph.
@@ 3.2.7 Step 7: Reviewing to check for inconsistency
After establishing the level partition, conical shape, and ISM model, it’s important to examine the full system to detect and resolve any discrepancies. This step verifies the system is consistent and functions smoothly without conflicting factors. The generated ISM model is reviewed for conceptual coherence, and any necessary revisions are implemented.
@@ 3.3 MICMAC analysis
MICMAC (multiplication of impact matrix applied to classification) study seeks to create a graph for categorizing various barriers. MICMAC algorithm is based on the multiplication characteristics of matrices [75].
To depict the interrelationships of a system’s components, the ISM framework requires the construction of a conical matrix. For each element, the driving power and dependence power are calculated by adding up the values in the rows and columns, respectively. These elements are further classified according to the effectiveness of their driving and dependent power in the MICMAC analysis. Factors are grouped into four categories: autonomous, dependent, linkage, and independent [76].
@@
    Group III (Linkage factors): Consists of linkage factors possessing robust propelling force and reliability.
@@
    Group IV (Independent factors): Involves factors that significantly impact the outcomes of sustainable energy initiatives but are not markedly influenced by other variables. These factors can be adeptly utilized to manipulate other elements, thereby contributing to the preservation of sustainable energy objectives.
@@ 3.4 DEMATEL
The DEMATEL method may transform the complex web of interrelations between barriers into a clear structural model of the system by categorizing them into two groups: causes and effects. It’s a practical method for ranking the relative importance of different aspects of a complex system to inform strategic decisions over the long run and identify potential development areas [77]. It is extensively implemented across numerous domains, such as SCM, traceability, smart cities, healthcare, and consumer behavior [78]. The following is a detailed description of how DEMATEL works [13] [27].
@@ 3.4.1 Step 1: Computation of average matrix
Everyone on the expert panel that filled out the survey was asked to provide an integer value between 0 and 4 to represent the degree of influence they saw between any two barriers, utilizing five different integer scales as “very high influence (4)”, “high influence (3)”, “moderate influence (2)”, “low influence (1)”, and “no influence (0)”. A respondent’s estimate of how much variable $i$ influences variable $j$ is denoted by the $x_{ij}$ symbol. For $i=j$, 0 is assigned to each diagonal element.
For every respondent, we may generate a non-negative $n\times n$ matrix using the formula $X^{k}=[x_{ij}^{k}]$, where $k$ means that there were a total of responders with $1\leq k\leq H$ and factor’s total number is denoted by $n$. In this case, the $H$ responders’ matrices are denoted as $X^{1},X^{2},X^{3},...,X^{H}$.
@@ 3.4.4 Step 4: Calculation of cause and effect values
$R$ and $C$ values, representing the sum of the values in rows and columns, are now used in the calculations. Where,
@@
The transpose value of the $j^{th}$ column, denoted by $\left[C_{j}\right]^{{}^{\prime}}$, the sum of the values of the $j^{th}$ column, $C_{j}$, shows the effect of the $j^{th}$ attribute on the other attributes. The influence of the $i^{th}$ row, denoted by $r_{i}$, indicates the cause of the $i^{th}$ attribute on the other attributes. The sum $(R_{i}+C_{i})$ represents the cumulative effects experienced and exerted by factor $i$. It signifies the degree of importance of element $i$ within the overall system. The difference $(R_{i}-C_{i})$ indicates the net effects that factor $i$ contributes to the system. If the difference between $D_{i}$ and $R_{i}$ is positive, the influence factor $i$ may be seen as a net cause. Conversely, if the difference between $D_{i}$ and $R_{i}$ is negative, factor $i$ can be regarded as a net receiver.
@@ 3.4.5 Step 5: Calculation of the threshold value
The determination of the threshold value is necessary in order to construct the interrelationship map. The interrelationship map only displays equations whose penetration matrix values are greater than the threshold value. The mean of the elements inside the entire penetration matrix is computed using the Equation 6.
@@ 3.4.6 Step 6: Drawing interrelationship map
The interrelationship map is created to comprehensively analyse the final solution, considering the values of $R_{i}+C_{i}$ and $R_{i}-C_{i}$ for each attribute in relation to the threshold value.
@@ 3.4.7 Step 7: Calculation of the final ranking of attributes
The interrelationship map derives a potential framework and hierarchy for the factors. This is achieved by organizing the values of $R_{i}+C_{i}$ in a descending order.
@@ 4.1 ISM-MICMAC application
The ISM and MICMAC techniques are strong ways to create a hierarchy and reveal contextual interactions between IoT adoption barriers, especially regarding CSC. The principal aim of utilizing MICMAC is to evaluate the importance of dimensions that affect the system. With the use of this methodology, researchers can examine the complex interrelationships and dependencies between different components, identifying important factors that have a major impact on the system as a whole. The ISM-MICMAC technique has two main phases, which have been applied in various settings such as supply chains, industry, and healthcare. A hierarchical structure of chosen IoT adoption barriers is first constructed using ISM. The relative importance and hierarchy of each barrier are then determined by methodically examining cause-and-effect interactions. After the dimensions are established, professionals evaluate the correlations, which results in the development of a binary relation matrix that is then converted into a reachability matrix. The hierarchy of the barriers is displayed in this matrix, highlighting their key and consequential functions within the framework of CSC. The MICMAC technique is used in the next step to further explore the relationships and interactions between the various IoT adoption barriers that have been identified. The emphasis turns to understanding the dynamic behavior of barriers and classifying them according to their reliance and driving force. Finding driving barriers with high influence but low dependence and dependent barriers with high influence but low driving power is made easier using MICMAC analysis. With this categorization, the role of each barrier in the bigger picture and the vital links that exist between them can be better understood.
@@ 4.1.1 Structural self‑interaction matrix (SSIM) development
Experts’ judgments on the relationships between 13 identified barriers are used to inform the SSIM’s development, which involves using four symbols to depict these interactions in a pairwise fashion described in subsubsection 3.2.1. In Table 4, SSIM depicts element intrinsic linkages and internal dynamics.
@@ 13 Cold chain network
The symbols V, A, X, and O are used to represent the specific sort of connection between the two components $(i,j)$ being examined. The notation “V” indicates that element $i$ will exert influence on element $j$. A: signifies that element $i$ will be impacted by element $j$.The symbol “X” denotes a bidirectional link, indicating that elements $i$ and $j$ will mutually impact each other. There is no relationship or connection between the elements of the symbol “O”.
@@ 4.1.2 Reachability matrix formation
Four symbols of Table 4 are converted to 0s and 1s to implement SSIM in the first reachability matrix using rules described in subsubsection 3.2.2. A square matrix is prepared in Table 6 with binary values that show whether parts influence one other. The observed barriers are interconnected and have complicated relationships, which may be better understood using the reachability matrix.
@@ 4.1.3 Final reachability matrix formation
Transitivity is used to create the final reachability matrix in Table 7. Following the completion of the ISM technique’s analytical phase, the final reachability matrix is produced, which contains the discovered hierarchical linkages among system parts. The final reachability matrix determines the “driving and dependency power”, providing insight into the system’s dynamics. Analyzing “driving and dependency power,” Table 7 classifies parameters into “dependent, autonomous, linkage, and independent” categories. The barriers that display significant driving force and dependence are categorized as belonging to the linking group; none of them are classed as autonomous, non-dependent, or dependent.
@@ 4.1.4 Level partition
Barriers are broken down into three distinct sets: the reachability set, which contains the factor and all other factors it influences; the antecedent set, which contains the factor and all other factors that influence it; and the intersubsection set, which contains all factors in both of these sets.
Table 8 provides an overview of the final reachability matrix’s detailed reachability set, antecedent set, and inter subsection set of IoT implementation barrier in the CSC. All barriers have been categorized into levels.
@@ 4.1.5 Conical matrix
The conical matrix in Table 9 displays transitivity interactions between adoption barriers. Identifying indirect linkages between components shows the domino impact of system effects. A more complete picture of the system’s dynamics and structure is achieved with the use of this matrix, which assists in understanding the wider effect of changing one element on other elements. The conical shape of the final reachability matrix is devised in Figure 3.
@@
The hierarchical relationships among different system parts are better understood using a conical matrix (Table 9), which is a tool for creating conical structures. This conical form clarifies the system’s architecture by visually depicting the hierarchical connections.
@@ 4.1.6 MICMAC analysis
MICMAC analysis is a useful tool for determining the drive and dependence power of barriers. This analysis is conducted with the use of barriers in order to divide the structure into separate groups.
@@
Autonomous barriers: These barriers are located in the first quadrant. They aren’t very strong at either driving or being dependent.
Dependent barriers: These barriers are shown inside Quadrant-II. The driving power of the subject is weak, whereas their dependence power is strong.
Linkage barriers: These are inside Quadrant-III and exhibit strong driving and dependent power. In general, the instability of these barriers arises from the interconnectedness of their actions since any action done on one barrier might have consequences on others and could affect the barrier itself.
@@ 4.1.7 ISM model
This digraph of Fig 5 shows the barriers to preventing IoT adoption. It was made to identify the hierarchy levels associated with implementing IoT in cold SC. The conceptual coherence of the ISM model produced is checked, and any required adjustments are made.
@@ 4.2.1 Average matrix formation
Each expert on the panel filled out a survey with an integer value between 0 and 4 to indicate the level of impact they noticed between every given pair of obstacles. Using the matrix format of the survey data, an average matrix is determined using Equation 1 in Table 10.
@@ 4.2.2 Normalized direct relation matrix formation
Using Equation 2, the normalized direct relation matrix is formulated from the average matrix in Table 11. Maximum summation is 32.333, which functions as a divisor, and every cell value from the average matrix is divided by this value.
@@ 4.2.3 Total relation matrix formation
The total relation matrix is formulated using Equation 3 in Table 12. The values in the matrix indicate the degree of relation between pairs of factors. A higher value suggests a stronger relationship.
@@ 4.2.4 Identification of cause and effect
Table 13 provides information on the degree of influence of various factors related to IoT adoption in cold SC barriers. Each row represents a specific factor, while each column provides different metrics or totals. Row total and column total are calculated using Equation 4 and Equation 5. Causes have $R_{i}-C_{i}$ values greater than 0, whereas effects have smaller $R_{i}-C_{i}$ values.
@@
Regularity Compliance 2.050 0.178 2.228 1.872 Cause
@@ 4.2.5 Developing interrelationship map
A digraph illustrating the influential relationship between the barriers is shown in Fig 6. The threshold value calculated using Equation 6 is 0.099. Once the threshold value has been determined, the interrelationship map is produced. In this map, values above the threshold reflect the influence of barrier $i$ on barrier $j$. In contrast, when the value is lower than the threshold value, barrier $i$ is influenced by barrier $j$. The arrows serve to demonstrate the influence, demonstrating the important link among barriers. Values that are positioned above the neutral horizontal line are categorized as belonging to the cause group, while values that are positioned below the neutral horizontal line are categorized as belonging to the effect group.
@@ 5 Discussions
The numerous and interconnected barriers to extensive IoT implementation in the cold SC are analyzed in this thesis. The data from the survey was double-checked to make sure it was accurate. The Cronbach’s alpha test is used to ensure that survey results are acceptable and valid. Reliable measurement is ensured by this coefficient, which evaluates the internal consistency of the survey questions. After it has been confirmed that the alpha value is above the acceptable range, the subsequent stage of the process is initiated.
@@
In order to ascertain, The ISM-MICMAC-DEMATEL method is constructed to determine important barriers; this section is divided into three subsections: subsection 4.1 for ISM, subsection 4.1.6 for MICMAC and the other subsection 4.2 for DEMATEL.
@@
The ISM model was created to determine the hierarchy levels of IoTBs so CSC practitioners could see their dependence links and focus on the biggest adoption barriers to IoT deployment. This research identified five hierarchical layers to explain the linkages among chosen IoTBs. The implementation barrier at higher levels exhibits a greater susceptibility to the effect of other factors, whilst components situated at lower levels strongly influence other factors.
@@
Variables that are classified as high levels, namely Levels 1 and 2, are prone to being influenced by other variables. We identified acceptance, adoption, and reliability issues at the top of the hierarchy, driven by the four hierarchies below it. At the second level, we identified two barriers: high implementation cost and ROI uncertainty, caused by three hierarchical layers below it. On the third tier, there are six interconnected barriers that have mutual effects. These barriers are a lack of skilled workforce, data security and privacy, technological infrastructure, scalability challenges, cultural and ethical considerations, data integration and management, and extreme environment. Conversely, components categorized as low levels, namely Levels 4 and 5, can influence other factors more. IoT adoption barriers in level 4 are cold chain networks and level 5 is regulatory compliance. These factors serve as the foundational elements within the broader network of relationships.
@@
From MICMAC analysis, independent barriers are cold chain network and regulatory compliance; dependent barriers are ROI uncertainty and high implementation cost. The other seven are linkage barriers, and no autonomous barriers exist. Among the many forms of barriers, linkage, and independent barriers exhibit significant driving forces, making them crucial barriers.
@@
The interdependence of barriers is depicted on the interrelationship map. Regularity compliance, lack of skilled workforce, data security and privacy, technological infrastructure, scalability challenges, cultural and ethical considerations, data integration and management issues, extreme environment issues, and cold chain networks are in the cause group. High implementation cost, acceptance and adoption, reliability issues, and ROI uncertainty are in the effect group. The regularity compliance and cold chain network have been identified as the most important causes, respectively, with respective values of 1.8719 and 1.2308. These two barriers drive more in the larger context. Thus, improving its performance may promote system development. Reliability issues, acceptance and adoption, and high implementation costs are the most important effects influenced by other barriers. These effects have been quantified with values of -1.7808, -1.4118, and -94707, respectively. In accordance with the final ranking of attributes, the most important criteria are technological infrastructure, data security and privacy, and data integration and management. The criterion of extreme environment is of the least importance.

### v4

@@ 3.2.2 Step 2: Developing reachability matrix
The SSIM generates a reachability matrix; four symbols are converted to 0s and 1s to implement the SSIM in the initial reachability matrix. matrix.
@@ 3.2.4 Step 4: Developing level partition
Level partitioning refers to categorizing or classifying distinct levels inside the information system. This involves dividing the system into levels or hierarchies to better understand its structure at various levels. The resultant final reachability matrix is segmented into several hierarchical levels.
@@ 3.2.5 Step 5: Developing conical form
The conical form describes structuring or organizing an information system in an arrangement that resembles a conical shape, indicating a hierarchical structure. Subsequently, a conical form of the reachability matrix was created.
@@ 3.2.6 Step 6: Developing ISM model
A directed graph was constructed using a reachability matrix to eliminate transitive linkages. By exchanging the nodes of the variables for statements, they converted into an ISM from the resultant digraph.
@@ 3.2.7 Step 7: Reviewing to check for inconsistency
After establishing the level partition, conical shape, and ISM model, it’s important to examine the full system to detect and resolve discrepancies. This step verifies that the system was consistent and functions smoothly without conflicting factors. The generated ISM model is reviewed for conceptual coherence, and any necessary revisions were implemented.
@@ 3.3 MICMAC analysis
The multiplication of impact matrix applied to classification (MICMAC) study seeks to create a graph to categorize various barriers. The MICMAC algorithm is based on the multiplication characteristics of matrices [86].
The ISM framework requires the construction of a conical matrix to depict the interrelationships of a system’s components. For each element, the driving and dependence powers were calculated by adding the values in the rows and columns, respectively. These elements were further classified according to the effectiveness of their driving and dependent power in the MICMAC analysis. Factors were grouped into four categories: autonomous, dependent, linkage, and independent [87].
@@
    Group III (Linkage factors): Consists of linkage factors possessing a robust propelling force and reliability.
@@
    Group IV (Independent factors): Involves factors that significantly impact the outcomes of sustainable energy initiatives but are not significantly influenced by other variables. These factors can be utilized to manipulate other elements, thereby contributing to the preservation of sustainable energy objectives.
@@ 3.4 DEMATEL
The DEMATEL method transforms the complex web of interrelations between barriers into a clear structural model of the system by categorizing them into two groups: causes and effects. This is a practical method for ranking the relative importance of different aspects of a complex system to inform strategic decisions over the long run and identify potential development areas [88]. It is extensively implemented across numerous domains, such as SCM, traceability, smart cities, healthcare, and consumer behavior [89]. The following is a detailed description of how DEMATEL works [15] [37].
@@ 3.4.1 Step 1: Computation of average matrix
The expert panel that filled out the survey was asked to provide an integer value between 0 and 4 to represent the degree of influence they saw between any two barriers, using five different integer scales as “very high influence (4)”, “high influence (3)”, “moderate influence (2)”, “low influence (1)”, and “no influence (0)”. A respondent’s estimate of how much variable $i$ influences variable $j$ is denoted by the $x_{ij}$ symbol. For $i=j$, 0 is assigned to each diagonal element.
For every respondent, we may generate a non-negative $n\times n$ matrix using the formula $X^{k}=[x_{ij}^{k}]$, where $k$ means that there were a total of responders with $1\leq k\leq H$ and the total number of factors is denoted by $n$. In this case, the $H$ responder matrices are denoted by $X^{1},X^{2},X^{3},...,X^{H}$.
@@ 3.4.4 Step 4: Calculation of cause and effect values
$R$ and $C$ values, representing the sum of the values in the rows and columns, are now used in the calculations. Where,
@@
The transpose value of the $j^{th}$ column, denoted by $\left[C_{j}\right]^{{}^{\prime}}$, and the sum of the values of the $j^{th}$ column, $C_{j}$, shows the effect of the $j^{th}$ attribute on the other attributes. The influence of the $i^{th}$ row, denoted by $r_{i}$, indicates the cause of the $i^{th}$ attribute on other attributes. The sum $(R_{i}+C_{i})$ represents the cumulative effects experienced and exerted by factor $i$. This signifies the degree of importance of element $i$ in the overall system. The difference $(R_{i}-C_{i})$ indicates the net effects of factor $i$ on the system. If the difference between $D_{i}$ and $R_{i}$ is positive, the influence factor $i$ may be considered as a net cause. Conversely, if the difference between $D_{i}$ and $R_{i}$ is negative, then factor $i$ can be regarded as a net receiver.
@@ 3.4.5 Step 5: Calculation of the threshold value
Determination of the threshold value is necessary to construct an interrelationship map. The interrelationship map only displays equations with penetration matrix values greater than the threshold value. The mean of the elements inside the entire penetration matrix is computed using Equation 6.
@@ 3.4.6 Step 6: Drawing interrelationship map
An interrelationship map is created to comprehensively analyze the final solution, considering the values of $R_{i}+C_{i}$ and $R_{i}-C_{i}$ for each attribute in relation to the threshold value.
@@ 3.4.7 Step 7: Calculation of the final ranking of attributes
The interrelationship map derives a potential framework and hierarchy of these factors. This is achieved by organizing the values of $R_{i}+C_{i}$ in descending order.
@@ 4.1 ISM-MICMAC application
ISM and MICMAC techniques are strong ways to create a hierarchy and reveal contextual interactions between IoT adoption barriers, especially regarding the CSC. The principal aim of utilizing MICMAC is to evaluate the importance of the dimensions that affect a system. Using this methodology, researchers can examine the complex interrelationships and dependencies between different components and identify important factors that have a major impact on the system as a whole. The ISM-MICMAC technique has two main phases that have been applied in various settings such as supply chains, industry, and healthcare. A hierarchical structure of the chosen IoT adoption barriers was first constructed using the ISM. The relative importance and hierarchy of each barrier were then determined by methodically examining cause-and-effect interactions. After the dimensions are established, professionals evaluate the correlations, resulting in the development of a binary relation matrix that is then converted into a reachability matrix. The hierarchy of the barriers is displayed in this matrix, highlighting their key and consequential functions within the CSC framework. The MICMAC technique is used in the next step to further explore the relationships and interactions among the various IoT adoption barriers that have been identified. The emphasis is on understanding the dynamic behavior of barriers and classifying them according to their reliance and driving forces. Identifying driving barriers with high influence but low dependence and dependent barriers with high influence but low driving power is made easier using MICMAC analysis. With this categorization, the role of each barrier in the bigger picture and the vital links between them can be better understood.
@@ 4.1.1 Structural self‑interaction matrix (SSIM) development
Experts’ judgments on the relationships between the 13 identified barriers are used to inform the SSIM’s development, which involves using four symbols to depict these interactions in a pairwise fashion, as described in subsubsection 3.2.1. In Table 4, SSIM depicts the element intrinsic linkages and internal dynamics.
@@ 13 Cold chain network
The symbols V, A, X, and O represent the specific sort of connection between the two components $(i,j)$ being examined. The notation “V” indicates that element $i$ will exert influence on element $j$. “A” signifies that element $i$ will be impacted by element $j$.The symbol “X” denotes a bidirectional link, indicating that elements $i$ and $j$ will mutually impact each other. There is no relationship or connection between the elements of the symbol”O”.
@@ 4.1.2 Reachability matrix formation
The four symbols in Table 4 are converted to 0s and 1s to implement SSIM in the first reachability matrix using the rules described in subsubsection 3.2.2. A square matrix is prepared in Table 6 with binary values that show whether the parts influence one other. The observed barriers are interconnected and have complicated relationships, which can be better understood using the reachability matrix.
@@ 4.1.3 Final reachability matrix formation
Transitivity was used to create the final reachability matrix, as shown in Table 7. Following the completion of the analytical phase of the ISM technique, the final reachability matrix, which contains the discovered hierarchical linkages among the system parts, is produced. The final reachability matrix determines the “driving and dependency power”, providing insight into the dynamics of the system. Analyzing “driving and dependency power,” Table 7 classifies parameters into “dependent, autonomous, linkage, and independent” categories. The barriers that display a significant driving force and dependence are categorized as belonging to the linking group; none of them are classified as autonomous, non-dependent, or dependent.
@@ 4.1.4 Level partition
Barriers are broken down into three distinct sets: the reachability set, which contains the factor and all other factors it influences; the antecedent set, which contains the factor and all other factors that influence it; and the intersubsection set, which contains all the factors in both sets.
Table 8 provides an overview of the final reachability matrix’s detailed reachability set, antecedent set, and inter-subsection set of IoT implementation barriers in CSC. All the barriers were categorized into levels.
@@ 4.1.5 Conical matrix
The conical matrix in Table 9 displays the transitivity interactions between adoption barriers. Identifying indirect linkages between components reveals the domino impact of system effects. A more complete picture of the system dynamics and structure is achieved using this matrix, which assists in understanding the wider effect of changing one element on other elements. The conical shape of the final reachability matrix is illustrated in Figure 3.
@@
The hierarchical relationships among the different system parts are better understood using a conical matrix (Table 9), which is a tool for creating conical structures. This conical form clarifies the architecture of the system by visually depicting the hierarchical connections.
@@ 4.1.6 MICMAC analysis
MICMAC analysis is a useful tool for determining the drive and dependence power of barriers. This analysis was conducted using barriers to divide the structure into separate groups.
@@
Autonomous barriers: These barriers were located in the first quadrant. They were not very strong in driving or being dependent.
Dependent barriers: These barriers are shown in Quadrant-II. The driving power of the subject was weak, whereas the dependence power was strong.
Linkage barriers: These are inside Quadrant-III and exhibit strong driving and dependent power. In general, the instability of these barriers arises from the interconnectedness of their actions since any action done on one barrier might have consequences for others and could affect the barrier itself.
@@ 4.1.7 ISM model
The digraph in Fig 5 shows the barriers preventing IoT adoption. It was used to identify the hierarchy levels associated with the implementation of the IoT in cold SC. The conceptual coherence of the produced ISM model was checked, and any required adjustments were made.
@@ 4.2.1 Average matrix formation
Each expert on the panel completed a survey with an integer value between 0 and 4 to indicate the level of impact they noticed between every given pair of obstacles. Using the matrix format of the survey data, the average matrix is determined using Equation 1 in Table 10.
@@ 4.2.2 Normalized direct relation matrix formation
Using Equation 2, the normalized direct relation matrix is formulated from the average matrix in Table 11. The maximum summation was 32.333, which functions as a divisor, and every cell value from the average matrix was divided by this value.
@@ 4.2.3 Total relation matrix formation
The total relation matrix is formulated using Equation 3 in Table 12. The values in the matrix indicate the degree of relation between the pairs of factors. A higher value suggests a stronger relationship.
@@ 4.2.4 Identification of cause and effect
Table 13 provides information on the degree of influence of various factors related to IoT adoption on cold SC barriers. Each row represents a specific factor, and each column provides different metrics or totals. The row total and column total were calculated using Equation 4 and Equation 5. Causes have $R_{i}-C_{i}$ values greater than 0, whereas the effects have smaller $R_{i}-C_{i}$ values.
@@
Regulatory Compliance 2.050 0.178 2.228 1.872 Cause
@@ 4.2.5 Developing interrelationship map
A digraph illustrating the influential relationship between the barriers is shown in Fig 6, where the threshold value calculated using Equation 6 is 0.099. Once the threshold value is determined, an interrelationship map is produced. In this map, the values above the threshold reflect the influence of barrier $i$ on barrier $j$. By contrast, when the value is lower than the threshold value, barrier $i$ is influenced by barrier $j$. The arrows demonstrate the influence, demonstrating an important link among the barriers. Values positioned above the neutral horizontal line are categorized as belonging to the cause group, while values positioned below the neutral horizontal line are categorized as belonging to the effect group.
@@ 5 Discussions
Numerous interconnected barriers to extensive IoT implementation in cold SC were analyzed in this study. Data from the survey were double-checked to ensure accuracy. The Cronbach’s alpha test was used to ensure that the survey results were acceptable and valid. A reliable measurement is ensured by this coefficient, which evaluates the internal consistency of the survey questions. After it was confirmed that the alpha value was above the acceptable range, the subsequent stage of the process was initiated.
@@
To ascertain this, The ISM-MICMAC-DEMATEL method is constructed to determine important barriers, which are divided into three subsections: subsection 4.1 for ISM, subsection 4.1.6 for MICMAC and the other subsection 4.2 for DEMATEL.
@@
The ISM model was created to determine the hierarchy levels of IoTBs so that CSC practitioners could see their dependence links and focus on the largest adoption barriers to IoT deployment. This study identified five hierarchical layers to explain the linkages between the chosen IoTBs. The implementation barrier at higher levels exhibits greater susceptibility to the effect of other factors, while components situated at lower levels strongly influence other factors.
@@
Variables classified as high levels, namely Levels 1 and 2, are prone to being influenced by other variables. We identified acceptance, adoption, and reliability issues at the top of the hierarchy, driven by the four hierarchies below. At the second level, we identified two barriers: high implementation cost and ROI uncertainty caused by the three hierarchical layers below it. In the third tier, six interconnected barriers have mutual effects. These barriers include a lack of skilled workforce, data security and privacy, technological infrastructure, scalability challenges, cultural and ethical considerations, data integration and management, and extreme environments. Conversely, components categorized as low levels, namely Levels 4 and 5, can influence the other factors more. The IoT adoption barriers in Level 4 are cold chain networks and Level 5 is regulatory compliance. These factors serve as foundational elements in a broader network of relationships.
The analysis of barriers identifies regulatory compliance as an important factor that hinders the implementation of the IoT. The implementation of government regulations that promote the adoption of IoT will result in a significant transformation of the supply chain system. Infrastructure, environmental factors, a trained workforce, cultural-ethical considerations, scalability, and other factors influence the cold chain network. Most of the supply chain stages are apprehensive about implementing IoT, as is typical with every new technology. An efficient technological infrastructure is necessary for the successful deployment of the IoT across supply chains, which involves a vast network of linked devices and systems. Adoption of the IoT depends on this infrastructure, which allows integration, data transfer, and communication among components in the supply chain ecosystem. Industries may face challenges with scalability and performance if the infrastructure is insufficient for the IoT. To oversee such a vast and complex system, experts with the necessary level of expertise are required.  Reducing the challenges associated with IoT adoption requires secure network cloud data. Preventing spyware and attacks is crucial since the number of connected devices and sensitive data is growing. As the nature of threats evolves, more sophisticated security measures and solutions designed for IoT systems may be required. A massive investment is necessary to hire experts, educate current employees, and build infrastructure. Industries are hesitant to implement IoT due to concerns about the potential lack of return on investment. In order to feel certain about the ROI, small companies need to be presented with a handful of positive examples.
@@
According to the MICMAC analysis, independent barriers are cold chain networks and regulatory compliance; dependent barriers are ROI uncertainty and high implementation costs. The other seven were linkage barriers, and no autonomous barriers existed. Among the many types of barriers, linkage and independent barriers exhibit significant driving forces, making them crucial barriers.
@@
The interdependence of the barriers is depicted in the interrelationship map. Regulatory compliance, lack of skilled workforce, data security and privacy, technological infrastructure, scalability challenges, cultural and ethical considerations, data integration and management issues, extreme environment issues, and cold chain networks are in the cause group. High implementation cost, acceptance and adoption, reliability issues, and ROI uncertainty are in the effect group. Regulatory compliance and cold chain networks were identified as the most important causes, with respective values of 1.8719 and 1.2308. These two barriers drive more in a larger context; thus, improving performance may promote system development. Reliability issues, acceptance and adoption, and high implementation costs are the most important effects influenced by other barriers. These effects were quantified with values of -1.7808, -1.4118, and -94707, respectively. According to the final ranking of attributes, the most important criteria are technological infrastructure, data security and privacy, and data integration and management. The criterion for extreme environments is of the least importance.
To overcome the barriers associated with data integration and management, it is essential to implement standardized software that enables communication between devices on the same system. Priority should be given to privacy and security during IoT implementation, and it is difficult to develop and investigate a gateway that connects devices using multiple protocols. Faculty and industry must collaborate to resolve this challenge. Significant investment and research are needed for solving security and data integration challenges at the ground level.